# SegChange-R1:Augmented Reasoning for Remote Sensing Change Detection via Large Language Models

### v1

@@ 3.3 Mask Decoder
The mask decoder serves as the core component responsible for generating precise segmentation masks by effectively integrating multi-scale visual features with semantic guidance from textual descriptions. Drawing inspiration from recent advances in language-guided segmentation [2], our architecture employs a hierarchical design that combines cross-modal feature fusion with transformer-based spatial reasoning capabilities. As shown in Fig. 2, the decoder comprises three primary components: a description projector (D-Projector) that bridges the semantic gap between language and vision domains, a transformer decoder that performs spatial reasoning through learnable queries, and a mask prediction head that generates the final segmentation output.
@@
The D-Projector module, as shown in Fig. 2, first aggregates temporal text embeddings and projects them into the visual feature space. Then, it uses a cross-attention mechanism to achieve fine-grained alignment between text semantics and spatial visual features. Subsequently, the transformer decoder interacts with the fused visual features through self-attention and cross-attention operations using a set of learnable queries, enabling the model to capture the long-range spatial dependencies and contextual associations required for accurate segmentation. The final mask prediction head employs a multi-scale convolutional architecture combined with channel attention mechanisms to refine spatial correlations and generate high-quality segmentation masks. This design enables the model to effectively handle complex spatial reasoning tasks while maintaining computational efficiency, as validated by our experimental results on change detection benchmarks.
@@
Table 2: Comparison results between SegChange-R1 and all state-of-the-art CD methods. The best results are shown in bold, and the second-best results are underlined. Our method outperforms existing methods on the DSIFN-CD, CDD, and DVCD datasets.
4 Experimental Results
Datasets:
@@ 4 Experimental Results
In recent years, with the rapid development of drone aerial photography technology, its advantages of flexible deployment, high spatial resolution, and multi-angle imaging have opened up new possibilities for precise monitoring of urban environments. However, most change detection datasets are still based on traditional remote sensing platforms, lacking explicit modeling at the semantic level. They are constrained by low spatial resolution and fixed shooting angles, making it difficult to accurately capture the diversity and fine-grained structures of building detail changes in complex urban environments. To further enhance the model’s ability to understand semantic changes and explore the auxiliary role of textual information in building change detection tasks.
@@
This study constructed a new drone-based change detection dataset, DVCD (Drone View Change Detection). This dataset primarily consists of drone orthophotos collected from most urban areas in Guangdong Province, covering building changes from 2022 to 2024, and accurately reflects typical change patterns such as new construction, demolition, and expansion/renovation during the rapid development of cities. Specifically, we introduced fine-grained textual description information into the dataset to guide the model to focus on semantic-level change features, thereby enhancing its ability to recognize the dynamic evolution of buildings in complex scenarios. The DVCD dataset comprises 12,833 pairs of dual-temporal images, divided into a training set of 11,066 pairs, a validation set of 1,383 pairs, and a test set of 1,384 pairs.
@@
To further validate the algorithm’s generalization ability, this study also conducted comparative experiments on three publicly available building change detection benchmark datasets, with details as shown in Table. 1. The WHU-CD [16] dataset is specifically designed for building extraction and change detection tasks; the DSIFN-CD [49] dataset consists of large-scale high-resolution dual-time-phase images covering six cities in China; The CDD [25] dataset provides diverse image pairs with varying resolutions and seasonal changes. These datasets each have unique characteristics in terms of image size, sample quantity, temporal span, spatial variation, and geographical coverage, providing a reliable foundation for a comprehensive evaluation of algorithm performance.
Experimental Setup:
@@
All experiments in this study were conducted on servers equipped with NVIDIA A800 GPUs. The training process used the AdamW optimizer, with an initial learning rate of $10^{-4}$, a backbone network learning rate of $10^{-5}$, and a weight decay coefficient of $10^{-4}$. The total number of training epochs was set to 128, with a batch size of 16, and a StepLR learning rate scheduling strategy was employed, reducing the learning rate to 0.1 times its original value every 20 epochs. During the testing phase, the batch size is set to 1, and the change detection threshold is set to 0.5. The model performance is evaluated using three widely used metrics: F1-Score to measure the harmonic mean of precision and recall, Intersection over Union (IoU) to assess the overlap between predicted and ground truth regions, and Overall Accuracy (OA) to measure the overall classification accuracy.
Results:
@@
We compare our method with recent state-of-the-art change detection methods on four datasets. As shown in Table. 2, our proposed SegChange-R1 demonstrated better performance in all benchmark tests. Among the existing methods, CNN-based models (e.g., FC-Siam-conc and [18]) have limited effectiveness due to their limited global context modeling capabilities. Transformer-based approaches (e.g., ChangeFormer [3] and SwinSUNet [44]) improve performance by modeling long-range dependencies through a self-attentive mechanism. Notably, ChangeCLIP [51] introduces instruction-guided learning and utilizes visual language pre-training to achieve good results on WHU-CD (F1: 0.982) and CDD (F1: 0.979), highlighting the advantages of incorporating semantic guidance. In contrast, our SegChange-R1 further utilizes text-guided semantic understanding, which achieves the highest performance on DSIFN-CD (F1: 0.972) and CDD (F1: 0.988).Moreover, it achieves the highest accuracy on the challenging drone-view DVCD dataset, highlighting its effectiveness in complex, high-resolution scenarios.
@@
For the challenging DVCD dataset evaluation, we specifically selected the most advanced open-source and reproducible methods and strictly followed the training configurations specified in their original papers to ensure fair comparison. Notably, our SegChange-R1 demonstrated outstanding training efficiency, converging in just 64 epochs and significantly outperforming all competitors’ latest methods.
@@
Table 3: When comparing the performance of different backbone networks, the Swin Transformer method demonstrated the best performance.
@@
Table 4: Prompt Configuration Performance Comparison.
Table 5: Bird’s Eye View Module Effectiveness
@@ 5 Ablation Study
To comprehensively validate the design approach and quantify the individual contributions of each proposal component, we conducted a systematic ablation study on the DVCD dataset. This analysis provided key insights for architecture selection and demonstrated the necessity of each module in achieving optimal change detection performance.
Backbone:
@@
We investigated three representative backbone architectures to understand their fundamental capabilities in capturing multi-scale spatial features, which are critical for change detection tasks, as shown in Table. 3. Comparative analysis revealed the different computational paradigms and their impact on change detection effectiveness.
@@
While ResNet-50 provides a solid baseline, its CNN architecture limits global context modeling.HGNetV2 adopts a lightweight architecture design, achieving an optimal balance between computational efficiency and representation capabilities. The Swin Transformer backbone network demonstrates excellent performance through its hierarchical self-attention mechanism, which is naturally suited to the multi-scale characteristics of change detection tasks. This architectural advantage translates into the ability to capture both fine-grained local changes and global contextual relationships, resulting in more coherent and accurate change detection results.
Text Guidance Strategy Analysis:
@@
We systematically evaluate the gradual integration of textual semantic information to understand its role in enhancing change detection through multimodal reasoning capabilities, as shown in Table. 4. The baseline configuration without textual guidance relies solely on visual feature discrimination, which limits the model’s ability to integrate semantic context and domain-specific knowledge. This limitation is particularly evident in scenarios involving complex change categories, which require semantic understanding beyond pure visual pattern recognition.
@@
Static prompts introduce basic semantic awareness through fixed text templates, providing consistent semantic anchors to help the model associate visual patterns with semantic concepts. However, the rigid nature of static prompts limits their adaptability to diverse contextual scenarios and change types, resulting in limited improvement. Dynamic descriptions achieve context-aware text based on scene features and change patterns, enabling the model to generate more relevant semantic guidance and thereby enhance the alignment between visual features and semantic understanding. LLM-enhanced reasoning methods leverage the complex reasoning capabilities of large language models to provide fine-grained semantic analysis and contextual understanding. This integration enables the model to perform complex semantic reasoning, combining domain knowledge and logical reasoning to significantly improve the accuracy and robustness of change detection.
Bird’s Eye View Module:
@@
The BEV module addresses the fundamental challenge of perspective distortion in aerial imagery, which can significantly impact change detection accuracy, particularly for objects with varying heights and orientations.Without BEV correction, the model struggles with perspective-induced geometric inconsistencies that can lead to false positive detections and missed changes.This is particularly problematic in urban environments where building heights and viewing angles create complex perspective effects.
@@
The results in Table. 5 demonstrate that, simple geometric projection provides basic perspective normalization through fixed transformation matrices, offering initial correction for systematic perspective distortions. However, this approach lacks the flexibility to adapt to scene-specific geometric variations and complex topographical features.Learnable transformation introduces adaptive perspective correction through trainable parameters that can adjust to scene-specific characteristics. This approach enables the model to learn optimal perspective transformations for different types of scenes and change patterns, resulting in improved spatial consistency.
@@
Our multi-scale BEV approach integrates perspective information hierarchically across different feature resolutions, enabling comprehensive perspective correction that preserves both global geometric consistency and local spatial details. This hierarchical integration ensures that perspective correction benefits are propagated throughout the feature extraction process, leading to substantial improvements in both spatial consistency and edge preservation accuracy.
@@ 6 Conclusion
In this paper, we propose SegChange-R1, a novel remote sensing change detection method leveraging large language models (LLMs) for enhanced semantic reasoning. By integrating textual descriptions, our approach effectively focuses on regions of interest, improving the accuracy and efficiency of change detection. To address modality misalignment between multi-temporal images, we design a BEV space converter based on linear attention, which aligns features into a unified spatial representation, enhancing spatial consistency and global context modeling.To support the development and evaluation of vision-language based change detection, we also introduce the first drone-view building change detection dataset, DVCD, featuring 13,800 image pairs across diverse urban and rural scenes. Experimental results on four benchmark datasets demonstrate that SegChange-R1 achieves state-of-the-art performance, particularly in F1 score, IoU, and overall accuracy. Ablation studies further validate the effectiveness of both the text-guided reasoning and the BEV module in improving model performance.This work opens new directions for integrating LLMs into remote sensing tasks.

### v2

@@ 3.3 Mask Decoder
The mask decoder generates precise segmentation masks by integrating multi-scale visual features with semantic guidance from text. Inspired by language-guided segmentation [2], our architecture uses a hierarchical design that combines cross-modal feature fusion with transformer-based spatial reasoning. As shown in Fig. 2, the decoder includes: a description projector (D-Projector) that bridges the semantic gap between language and vision, a transformer decoder that performs spatial reasoning through learnable queries, and a mask prediction head that generates the segmentation output.
@@
The D-Projector, shown in Fig. 2, first aggregates temporal text embeddings and projects them into the visual feature space. Then, it uses cross-attention to achieve fine-grained alignment between text semantics and spatial visual features. Subsequently, the transformer decoder interacts with the fused visual features through self- and cross-attention using learnable queries, enabling the model to capture long-range spatial dependencies and contextual associations for accurate segmentation. The final mask prediction head employs a multi-scale convolutional architecture with channel attention to refine spatial correlations and generate high-quality segmentation masks. This design enables the model to handle complex spatial reasoning while maintaining efficiency, as validated by our experiments.
@@
Table 2: Quantitative comparison of SegChange-R1 against state-of-the-art change detection (CD) methods. Best and second-best results are indicated in bold and underlined, respectively. SegChange-R1 demonstrates superior performance on the DSIFN-CD, CDD, and DVCD datasets.
4 Experimental Setup and Results
Datasets
@@ 4 Experimental Setup and Results
The rapid development of drone aerial photography, with its flexible deployment, high resolution, and multi-angle imaging, has opened new possibilities for urban environment monitoring. However, most change detection datasets are based on traditional remote sensing platforms and lack semantic-level modeling. They are constrained by low resolution and fixed angles, making it difficult to capture diverse building details in complex urban environments. This motivates enhancing the model’s understanding of semantic changes and exploring the role of text in building change detection.
@@
We constructed DVCD (Drone View Change Detection), a new drone-based change detection dataset. This dataset consists of drone orthophotos collected from urban areas in Guangdong Province, covering building changes from 2022 to 2024. It reflects typical changes such as new construction, demolition, and expansion during rapid urban development. We introduced fine-grained textual descriptions to guide the model to focus on semantic change features, enhancing its ability to recognize building evolution in complex scenarios. The DVCD dataset comprises 12,833 image pairs, divided into a training set (11,066 pairs), a validation set (1,383 pairs), and a test set (1,384 pairs).
@@
To further validate the algorithm’s generalization ability, this study also conducted comparative experiments on three publicly available building change detection benchmark datasets, with details as shown in Table. 1. The WHU-CD [16] dataset is specifically designed for building extraction and change detection tasks; the DSIFN-CD [49] dataset consists of large-scale high-resolution dual-time-phase images covering six cities in China; The CDD [25] dataset provides diverse image pairs with varying resolutions and seasonal changes. These datasets have unique characteristics in image size, sample quantity, temporal span, spatial variation, and geographical coverage, providing a foundation for evaluating algorithm performance.
Experimental Setup
@@
All experiments were conducted on servers with NVIDIA A800 Graphics Processing Units (GPUs). Training used the AdamW optimizer with a learning rate of $10^{-4}$, a backbone network learning rate of $10^{-5}$, and a weight decay of $10^{-4}$. Training used 128 epochs with a batch size of 16 and a StepLR scheduler, reducing the learning rate by a factor of 0.1 every 20 epochs. During testing, the batch size was 1, and the change detection threshold was 0.5. Model performance was evaluated using F1 Score (F1-Score; harmonic mean of precision and recall), Intersection over Union (IoU), and Overall Accuracy (OA).
Results
@@
We compare our method with recent state-of-the-art change detection methods on four datasets. As shown in Table. 2, our proposed SegChange-R1 demonstrated better performance in all benchmark tests. Among the existing methods, CNN-based models (e.g., FC-Siam-conc and [18]) have limited effectiveness due to their limited global context modeling capabilities. Transformer-based approaches (e.g., ChangeFormer [3] and SwinSUNet [44]) improve performance by modeling long-range dependencies through a self-attentive mechanism. Notably, ChangeCLIP [51] introduces instruction-guided learning and utilizes visual language pre-training to achieve good results on WHU-CD (F1: 0.982) and CDD (F1: 0.979), highlighting the advantages of incorporating semantic guidance. In contrast, SegChange-R1 utilizes text-guided semantic understanding, achieving the highest performance on DSIFN-CD (F1: 0.972) and CDD (F1: 0.988), and achieves the highest accuracy on the drone-view DVCD dataset, highlighting its effectiveness in complex scenarios.
@@
For the DVCD dataset evaluation, we selected advanced open-source methods and followed their original training configurations for fair comparison. Notably, SegChange-R1 demonstrated high training efficiency, converging in 64 epochs and outperforming all competitors.
@@
Table 3: Performance comparison of backbone networks. The Swin Transformer backbone achieved the highest performance.
@@
Table 4: Performance comparison of different prompt configurations for [Task/Model].
@@
Table 5: Effectiveness of the bird’s-eye view (BEV) module in improving [Task/Metric].
@@ 5 Ablation Study
To validate the design and quantify the contribution of each component, we conducted an ablation study on the DVCD dataset. This analysis provided insights for architecture selection and demonstrated the necessity of each module for optimal change detection.
Backbone
@@
We investigated three representative backbone architectures to understand their fundamental capabilities in capturing multi-scale spatial features, which are critical for change detection tasks, as shown in Table. 3. The analysis revealed the impact of different computational paradigms on change detection effectiveness.
@@
While ResNet-50 provides a baseline, its CNN architecture limits global context modeling. HGNetV2 adopts a lightweight architecture, balancing computational efficiency and representation. The Swin Transformer demonstrates strong performance through its hierarchical self-attention, suited to the multi-scale characteristics of change detection. This enables the capture of both fine-grained local changes and global context, resulting in more coherent change detection.
Text Guidance Strategy Analysis
@@
We systematically evaluate the gradual integration of textual semantic information to understand its role in enhancing change detection through multimodal reasoning capabilities, as shown in Table. 4. The baseline without text relies solely on visual features, limiting the model’s ability to integrate semantic context and domain knowledge. This limitation is evident in complex change categories, which require semantic understanding beyond visual recognition.
@@
Static prompts introduce semantic awareness through fixed text, providing consistent anchors to help the model associate visual patterns with semantic concepts. However, static prompts limit adaptability to diverse contexts and change types, resulting in limited improvement. Dynamic descriptions achieve context-aware text based on scene features and change patterns, enabling more relevant semantic guidance and enhancing alignment between visual features and semantics. LLM-enhanced reasoning leverages the reasoning capabilities of large language models to provide fine-grained semantic analysis and context. This enables the model to perform semantic reasoning, combining domain knowledge and logic to improve change detection accuracy and robustness.
Bird’s Eye View Module
@@
The BEV module addresses perspective distortion in aerial imagery, which can impact change detection accuracy, particularly for objects with varying heights and orientations. Without BEV correction, the model struggles with perspective-induced geometric inconsistencies, leading to false positives and missed changes. This is problematic in urban environments where building heights and viewing angles create perspective effects.
@@
The results in Table. 5 demonstrate that, simple geometric projection provides basic perspective normalization through fixed transformation matrices, offering initial correction for systematic perspective distortions. However, this approach lacks the flexibility to adapt to scene-specific geometric variations and complex topography. Learnable transformation introduces adaptive perspective correction through trainable parameters that adjust to scene-specific characteristics, enabling the model to learn optimal transformations for different scenes and improve spatial consistency.
@@
Our multi-scale BEV integrates perspective information hierarchically across feature resolutions, enabling perspective correction that preserves both global geometric consistency and local spatial details. This ensures that perspective correction benefits are propagated throughout feature extraction, improving spatial consistency and edge preservation.
@@ 6 Conclusion
We propose SegChange-R1, a remote sensing change detection method leveraging LLMs for enhanced semantic reasoning. By integrating textual descriptions, our approach focuses on regions of interest, improving the accuracy and efficiency of change detection. To address modality misalignment, we designed a linear attention-based BEV converter, which aligns features into a unified spatial representation, enhancing spatial consistency and global context modeling. To support vision-language change detection, we introduce DVCD, a drone-view building change detection dataset featuring 13,800 image pairs across diverse urban and rural scenes. Experiments on four benchmark datasets demonstrate that SegChange-R1 achieves state-of-the-art performance in F1 score, IoU, and overall accuracy. Ablation studies validate the effectiveness of text-guided reasoning and the BEV module. This work opens new directions for integrating LLMs into remote sensing.